OpenStudy (christos):

Calculus help me differentiate this: d/dx (x^3- (7/x))^-2 My answer is -2(x^3-7/x)^-3(3x^2+7x^-2)

OpenStudy (michele_laino):

that's right!

OpenStudy (christos):

wolf alpha gives a different answer https://www.dropbox.com/s/xma83v325e10s8u/Screenshot%202015-12-02%2023.29.15.png?dl=0

OpenStudy (kainui):

I copy pasted your problem with an equals sign to your answer you wrote and Wolfram Alpha returned true! :P http://www.wolframalpha.com/input/?i=d%2Fdx%20(x%5E3-%20(7%2Fx))%5E-2%20%3D%20-2(x%5E3-7%2Fx)%5E-3(3x%5E2%2B7x%5E-2)&t=crmtb01

OpenStudy (christos):

Thank you guyssss ! :D:D

OpenStudy (christos):

many many kudos !

OpenStudy (michele_laino):

oops.. I have made a typo: \[\huge \begin{gathered} - 2{\left( {{x^3} - \frac{7}{x}} \right)^{ - 3}}\left( {3{x^2} + \frac{7}{{{x^2}}}} \right) = \hfill \\ \hfill \\ = - 2\frac{{\frac{{3{x^4} + 7}}{{{x^2}}}}}{{{{\left( {\frac{{{x^4} - 7}}{x}} \right)}^3}}} = \hfill \\ \hfill \\ = \frac{{ - 2x\left( {3{x^4} + 7} \right)}}{{{{\left( {{x^4} - 7} \right)}^3}}} \hfill \\ \end{gathered} \]